OpenStudy (anonymous):

Solve the following equation algebraically, if necessary round to two decimal places: abs(2x-3)+4=5x

OpenStudy (p0sitr0n):

|2x-3|=5x-4 2x-3=5x-4 or -2x+3=5x-4 1=3x or -7x=-1 x=1/3 or x=1/7

OpenStudy (p0sitr0n):

keep in mind that your abs() expression is symmetrical, so it can be either bigger or smaller than 0

OpenStudy (anonymous):

Thank you so much!! So if I had an absolute value of x^2+5x-7 I could treat it the same way??